What is the value of :


 


cot 76 cot 44 + cot 44 cot 16 - cot 76 cot16

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear karthik,

E = cot(76)cot(44) + cot(44)cot(16) - cot(76)cot(16)

First, use the formula, cot(A) = tan(90 - A). Then we have :

E = tan(14)tan(46) + tan(46)tan(74) - tan(14)tan(74)

Rearrange this so the larger numbers are first :
E = tan(46)tan(14) + tan(74)tan(46) - tan(74)tan(14)

Now we need these two formulae :

tan(A + B) = [tan(A) + tan(B)]/[1 - tan(A)tan(B)]
which rearranges to : tan(A)tan(B) = 1 - [tan(A) + tan(B)]/tan(A + B) .......... (1)
and
tan(A - B) = [tan(A) - tan(B)] / [1 + tan(A)tan(B)]
which rearranges to : tan(A)tan(B) = [tan(A) - tan(B)]/tan(A - B) - 1 ............ (2)

For the 1st expression of E, use (1) : tan(46)tan(14) = 1 - [tan(46) + tan(14)]/tan(60)
But tan(60) = √3, so,
tan(46)tan(14) = 1 - [tan(46 + tan(14)]/√3 = 1 - tan(46)/√3 - tan(14)/√3 ........... (3)

For the 2st expression of E, use (1) : tan(74)tan(46) = 1 - [tan(74) + tan(46)]/tan(120)
But tan(120) = -√3, so,
tan(74)tan(46) = 1 - [tan(74) + tan(46)]/(-√3) = 1 + tan(74)/√3 + tan(46)/√3 ..... (4)

For the 3st expression of E, use (2) : -tan(74)tan(14) = -{[tan(74) - tan(14)]/tan(60) - 1}
But tan(60) = √3, so,
-tan(74)tan(14) = -{[tan(74) - tan(14)]/√3 - 1} = -tan(74)/√3 + tan(14)/√3 + 1 .... (5)

Now add, (3) + (4) + (5), which gives the answer, E = 3.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi